# Integrating Neural Differential Forecasting with Safe Reinforcement Learning for Blood Glucose Regulation

### v1

@@ 1 Introduction
Type 1 Diabetes (T1D) is a chronic autoimmune disease affecting over 9.5 million people worldwide, characterized by the destruction of pancreatic $\beta$-cells and resulting in absolute insulin deficiency [alberti1998definition, davis2015cpeptide]. Individuals with T1D must carefully regulate blood glucose (BG) within a safe range (usually set between 70–180 mg/dL) [ada2022standards]. Failure to do so can lead to acute hypoglycemia or hyperglycemia and long-term vascular and neurological complications, or even coma and death [mol2004hypoglycaemia, mynatt2010ubiquitous, zhou2022openaps]. Despite progress in continuous glucose monitoring (CGM) and insulin pumps, fully autonomous insulin delivery remains challenging due to delayed glucose–insulin dynamics, inter-patient variability, and uncertainty in meal timing and size [Jacobs2025AIDChallenges].
@@
Our framework, termed TSODE, integrates three synergistic components: a Thompson Sampling policy for adaptive insulin selection, a NeuralODE forecaster for glucose trajectory prediction, and a conformal safety layer that enforces certified action filtering (see Fig 1 on the later page). Evaluated in the FDA-approved UVa/Padova Type 1 Diabetes Simulator [dallaman2014uva], TSODE achieves high time-in-range while substantially reducing hypoglycemia risk compared to RL-control and model-predictive baselines.
@@ 2 Related Work
Classical glucose control algorithms such as proportional integral derivative (PID) and model predictive control (MPC) have demonstrated strong performance in maintaining near-normoglycemia [bequette2013challenges, kovatchev2009artificial]. However, these approaches depend on explicit physiological models and manual parameter tuning, limiting their adaptability to patient variability and unannounced meals. Data-driven glucose forecasters based on deep learning [mirshekarian2022neural] can improve short-term prediction accuracy but cannot directly optimize sequential dosing decisions.
Offline reinforcement learning (RL) has recently emerged as a safer alternative for healthcare control problems where online exploration is risky [levine2020offline]. In the context of blood glucose regulation, several studies have shown that RL can enhance time-in-range (TIR) and personalization using only retrospective or simulated data. Examples include offline RL for safer insulin control [emerson2022offline], distributional and uncertainty-aware RL for titration [desman2025glucose], and deep actor–critic formulations for hybrid closed-loop systems [lee2023safe, zhu2024adaptive]. Complementary work has also combined glucose prediction with dose selection via deep Q-networks [symeonidis2025deep].
While these approaches improve glycemic regulation, they typically focus on maximizing average performance without explicitly constraining unsafe actions such as overdosing or rapid correction stacking. The proposed TSODE framework addresses this limitation by integrating predictive safety certification via a NeuralODE-based probabilistic forecaster and conformal calibration directly into a Thompson Sampling RL controller.
@@ 3.1 Overview
We design a discrete-action closed-loop controller that combines Thompson Sampling reinforcement with NeuralODE based safety supervision, as in Fig 1. The controller operates in discrete time. At each decision step, it observes recent glucose and contextual features, proposes an insulin action through a Thompson Sampling policy, and evaluates its safety using a probabilistic forecaster based on neural ordinary differential equations. Unsafe proposals are adaptively corrected through a constrained optimization step, while additional rule-based guardrails enforce hard physiological limits.
@@ 3.2 State and Action Space
At each decision step $t$, the controller forms a state vector $\mathbf{x}_{t-H+1:t}$ from recent glucose readings, insulin delivery, carbohydrate estimates, time-of-day, and risk metrics. The action space $\mathcal{A}$ is a discrete grid of insulin bolus increments within a physiologically safe range.
@@ 3.3 Thompson Sampling Policy
For each state–action pair $(s,a)$, the controller maintains online estimates of the mean and variance of the shaped reward using Welford’s algorithm. Action selection follows the standard Thompson Sampling rule, where each action’s expected reward is modeled as Gaussian, a sample $\tilde{r}_{a}$ is drawn from the posterior of each action, and the controller selects
@@ 3.4 Neural Ordinary Differential Equation Forecaster
To model continuous-time glucose–insulin dynamics, we employ a Neural Ordinary Differential Equation (NeuralODE) forecaster that predicts short-term glucose trajectories conditioned on recent physiological context and proposed insulin actions. At each control step, the model receives a sliding history window of standardized features (e.g., BG, IOB, COB, time-of-day, and risk indices) and encodes it through a gated recurrent unit (GRU) to produce a latent initial state $\mathbf{z}_{0}$.
@@ 3.5 Probabilistic Safety Filter
Let $\mu_{1:K}$ denote the forecasted BG trajectory (mg/dL) under a proposed bolus $u$, produced by injecting $u$ into the latest insulin feature before forecasting. Define
$W(\mu)=\sum_{k=1}^{K}w_{k}\,\mu_{k}\qquad S(\mu)=\frac{\mu_{K}-\mathrm{BG}_{t}}{K\,\Delta t}$ (3)
where exponential weights $w_{k}\propto\lambda^{k}$ emphasize near-term predictions and $\Delta t=3$ min. Here, $W(\mu)$ represents the weighted average of the predicted glucose trajectory and $S(\mu)$ estimates the average rate of glucose change over the prediction horizon. These statistics summarize serve as compact risk indicators for evaluating the safety of the proposed insulin action. The gate enforces
$\Pr\!\big(W(\mu)\geq L\big)$ $\geq 1-\alpha_{W}$ (4)
$\Pr\!\big(S(\mu)\geq-\gamma\big)$ $\geq 1-\alpha_{S}$ (5)
using $L=120$ mg/dL, $\gamma=0.3$ mg/dL min-1, $\alpha_{W}=0.018$, and $\alpha_{S}=0.002$ (total budget $\alpha=0.02$). Conformal quantiles $q_{W},q_{S}$ are estimated online from absolute residuals of $W$ and $S$, yielding tests
$W(\mu)-q_{W}\geq L\qquad S(\mu)+q_{S}\geq-\gamma$ (6)
Before sufficient calibration ($n<80$), we apply Cantelli bounds using horizon-wise residual standard deviations $\sigma_{h}$.
Safe Capping and Guardrails. If tests fail, a 1-D bisection over $u\in[0,u_{\mathrm{prop}}]$ returns the largest safe dose, else $u=0$. When BG is high and rising, the forecast gate is bypassed. Final delivery passes deterministic guardrails: glucose/trend thresholds, IOB caps, and a late-night cap.
@@
TIR% Time<<70% Mean BG TIR% Time<<70% Mean BG TIR% Time<<70% Mean BG TIR% Time<<70% Mean BG
@@ 4.1 Experimental Setting
Data Description. All experiments were performed using the SimGlucose environment, a Python interface built on the UVa/Padova Type 1 Diabetes simulator. The simulator models glucose–insulin dynamics for virtual adults with individualized metabolic parameters.
Experiment Setup. Each virtual adult patient was simulated for 44 consecutive days in the simulator. The first 30 days were used for policy warm-up and the last 14 for closed-loop evaluation. Every $\Delta t=3$ min, the controller observed a normalized feature vector $x_{t}$ (BG, IOB, COB, time-of-day, and risk indices) over a sliding window of $H=10$ steps (30 min). The discrete action set $\mathcal{A}=\{0,0.2,\ldots,3.0\}$ U represents physiologically safe bolus increments, with a short refractory period preventing rapid re-dosing. All controllers followed identical meal schedules (08:00, 12:30, 16:00, 19:00; 50–70–15–60 g) and were simulated with fixed random seeds for reproducibility.
Benchmark Algorithms. We compared TSODE with three standard baselines: (i) a rule-based Meal–Bolus controller using fixed ICR and basal rate, (ii) a classical PID loop maintaining BG near 120 mg/dL, and (iii) a hybrid Thompson Sampling + MPC (TSMPC) controller that adds an analytic safety layer but no neural forecasting.
@@ 4.2 Results
Fig. 3: Daily TIR by Patient
Figures 2 and 3 illustrate the closed-loop performance of the proposed TSODE controller. Figure 2 compares a representative 24-hour simulation against a conventional PID controller for the same virtual adult. While PID shows larger glucose excursions and periods of suboptimal regulation, TSODE maintains levels tightly within the target range of 70–180 mg/dL, achieving full time-in-range (TIR = 100%) with no hypoglycemia. Figure 3 extends this comparison across three adult subjects, where daily TIR values consistently cluster between 85% and 100%, reflecting stable glycemic control and strong generalization despite variability in metabolism and insulin sensitivity. Together, these results demonstrate that TSODE achieves safe, adaptive, and near-normoglycemic regulation superior to conventional PID control.
Table 1 summarizes mean performance across three virtual adult patients over 14 evaluation days. TSODE consistently achieves the highest TIR (87.9%) while reducing hypoglycemia to below 10%, outperforming PID and TSMPC baselines. While PID and TSMPC reach high TIR, they suffer from frequent overdosing and up to 26% low glucose time. By combining meal-aware adaptation with NeuralODE forecasting, TSODE anticipates downward trends and caps unsafe doses, enhancing both performance and safety in autonomous insulin delivery.
Model Transferability and Generalization. To evaluate cross-patient generalization, the TSODE controller was trained on adults#001 and #002 and tested on an unseen subject (adult#005). As summarized in Table 1, the Meal-Bolus baseline achieved only 21.2% TIR, while the PID controller improved TIR to 75.1% but incurred 24.8% time below 70 mg/dL. In contrast, TSODE maintained a high TIR of 88.9% with only 3.8% time << 70 mg/dL, achieving comparable control with substantially reduced hypoglycemia risk. These results highlight the controller’s ability to generalize across unseen patients while preserving safety through NeuralODE based probabilistic supervision.
@@ 5 Conclusion
We presented TSODE, a safety-aware controller that integrates TS reinforcement learning with NeuralODE based probabilistic forecasting for automated insulin delivery. TSODE adaptively regulates glucose while certifying each insulin action through conformal safety checks and guardrail constraints. In the FDA-approved UVa/Padova simulator, it achieved 87.9% time-in-range and only 6.1% time below 70 mg/dL, representing a 5–8 pp increase in TIR and a 65% reduction in hypoglycemia compared to PID and TSMPC baselines. These results demonstrate that probabilistic safety filtering substantially enhances both stability and safety of closed-loop glucose control.

### v2

@@ 1 Introduction
Type 1 Diabetes (T1D) is a chronic autoimmune disease affecting over 9.5 million people worldwide, characterized by the destruction of pancreatic $\beta$-cells and resulting in absolute insulin deficiency [alberti1998definition, davis2015cpeptide]. Individuals with T1D must carefully regulate blood glucose (BG) within a safe range (usually set between 70–180 mg/dL) [ada2022standards, zhou2021data, zhou2023hybrid]. Failure to do so can lead to acute hypoglycemia or hyperglycemia and long-term vascular and neurological complications, or even coma and death [mol2004hypoglycaemia, mynatt2010ubiquitous, zhou2022design]. Despite progress in continuous glucose monitoring (CGM) and insulin pumps, fully autonomous insulin delivery remains challenging due to delayed glucose–insulin dynamics, inter-patient variability, and uncertainty in meal timing and size [Jacobs2025AIDChallenges].
@@
Our framework, termed TSODE, integrates three synergistic components, including a Thompson Sampling policy for adaptive insulin selection, a NeuralODE forecaster for glucose trajectory prediction, and a conformal safety layer that enforces certified action filtering (see Fig. 1 on the later page). Evaluated in the FDA-approved UVa/Padova Type 1 Diabetes Simulator [dallaman2014uva], TSODE achieves higher time-in-range while reducing hypoglycemia risk more than 60% compared to RL-control and model-predictive baselines.
@@ 2 Related Work
Classical glucose control algorithms such as proportional integral derivative (PID) and model predictive control (MPC) have demonstrated strong performance in maintaining near-normoglycemia [bequette2013challenges, kovatchev2009artificial, bagheri2022towards]. However, these approaches depend on explicit physiological models and manual parameter tuning, limiting their adaptability to patient variability and unannounced meals. Data-driven glucose forecasters based on deep learning [mirshekarian2022neural, smith2023short] can improve short-term prediction accuracy but cannot directly optimize sequential dosing decisions.
Offline reinforcement learning (RL) has recently emerged as a safer alternative for healthcare control problems where online exploration is risky [levine2020offline]. In the context of blood glucose regulation, several studies have shown that RL can enhance time-in-range (TIR) and personalization using only retrospective or simulated data. Examples include offline RL for safe insulin control [emerson2022offline], distributional and uncertainty-aware RL for titration [desman2025glucose], and deep actor–critic formulations for hybrid closed-loop systems [lee2023safe, zhu2024adaptive].
While these approaches improve glycemic regulation, they typically focus on maximizing average performance without explicitly constraining unsafe actions such as overdosing or rapid correction stacking. Recent work [zhou2025knowsafe, symeonidis2025deep] addresses this limitation but relies on multiple DNN models, increasing system complexity. The proposed TSODE framework integrates predictive safety certification via a single NeuralODE-based probabilistic forecaster and conformal calibration directly into a Thompson Sampling RL controller.
@@ 3.1 Overview
We design a discrete-action closed-loop controller that combines Thompson Sampling reinforcement with NeuralODE forecasting-based safety supervision, as in Fig. 1. The controller operates in discrete time. At each decision step, it observes recent glucose and contextual features, proposes an insulin action through a Thompson Sampling policy, and evaluates its safety using a probabilistic forecaster based on neural ordinary differential equations. Unsafe proposals are adaptively corrected through a constrained optimization step, while additional rule-based guardrails enforce hard physiological limits.
@@ 3.2 State and Action Space
At each decision step $t$, the controller forms a state vector $\mathbf{x}_{t-H+1:t}$ from recent glucose readings, insulin delivery (IOB), carbohydrate estimates (COB), time-of-day, and risk metrics. The action space $\mathcal{A}$ is a discrete grid of insulin bolus increments within a physiologically safe range.
@@ 3.3 Thompson Sampling Policy
For each state–action pair $(s,a)$, the controller maintains numerically stable online estimates of the mean and variance of the shaped reward. Action selection follows the standard Thompson Sampling rule, where each action’s expected reward is modeled as Gaussian, a sample $\tilde{r}_{a}$ is drawn from the posterior of each action, and the controller selects
@@ 3.4 Neural Ordinary Differential Equation Forecaster
To model continuous-time glucose–insulin dynamics, we employ a NeuralODE forecaster that predicts short-term glucose trajectories conditioned on recent physiological context and proposed insulin actions. At each control step, the model receives a sliding history window of standardized features (BG, IOB, COB, time-of-day, and risk indices) and encodes it through a gated recurrent unit (GRU) to produce a latent initial state $\mathbf{z}_{0}$.
@@ 3.5 Probabilistic Safety Filter
Given a proposed bolus $u$, the NeuralODE forecaster predicts a glucose trajectory $\mu_{1:K}$, where $\mu_{k}$ denotes the predicted blood glucose (mg/dL) at future step $k$. Hypoglycemia risk [zhou2021data, zhou2023hybrid] is summarized using a weighted trajectory average and a horizon-level slope:
$W(\mu)$ $=\sum_{k=1}^{K}w_{k}\,\mu_{k},\qquad S(\mu)=\frac{\mu_{K}-\mathrm{BG}_{t}}{K\,\Delta t}$ (3)
where $K$ is the prediction horizon, $w_{k}$ are exponentially decaying weights emphasizing near-term predictions, $\mathrm{BG}_{t}$ is the current glucose value, and $\Delta t$ is the control interval.
Safety is enforced by requiring both statistics to exceed predefined thresholds with high probability:
$\Pr\!\big(W(\mu)\geq L,\;S(\mu)\geq-\gamma\big)\;\geq\;1-\alpha,$ (4)
These constraints enforce high-probability lower bounds on predicted glucose level and descent rate, respectively. Uncertainty is handled via conformal calibration using empirical residual quantiles. If tests fail, a 1-D bisection over $u\in[0,u_{\mathrm{prop}}]$ returns the largest safe dose, else $u=0$. When BG is high and rising, the forecast gate is bypassed. Final delivery passes deterministic guardrails: glucose/trend thresholds, IOB caps, and a late-night cap.
@@
TIR% Time<<70% Mean BG (mg/dL) TIR% Time<<70% Mean BG (mg/dL) TIR% Time<<70% Mean BG (mg/dL) TIR% Time<<70% Mean BG (mg/dL)
@@ 4.1 Experimental Setting
Data Description. All experiments were performed using a closed-loop testbed [zhou2022design] built on the UVa/Padova Type 1 Diabetes simulator. The simulator models glucose-insulin dynamics for virtual adults with individualized metabolic parameters. Real patient experiments are not conducted due to the safety-critical nature of insulin dosing.
Experiment Setup. Each virtual adult patient was simulated for 44 consecutive days in the simulation. The first 30 days were used for policy warm-up and the last 14 for closed-loop evaluation. Every $\Delta t=3$ min, the controller observed a normalized feature vector $x_{t}$ over a sliding window of $H=10$ steps (30 min). The discrete action set $\mathcal{A}=\{0,0.2,\ldots,3.0\}$ U represents physiologically safe bolus increments (in international units), with a short refractory period preventing rapid re-dosing. All controllers followed identical meal schedules (08:00, 12:30, 16:00, 19:00; 50-70-15-60 g) and were simulated with fixed random seeds for reproducibility. An open-source implementation is available at GitHub.
Benchmark Algorithms. We compared TSODE with three standard baselines: (i) a rule-based Meal-Bolus controller using fixed insulin-to-carbohydrate ratio (ICR) and basal rate, (ii) a classical PID loop maintaining BG near 120 mg/dL, and (iii) a hybrid MPC-based controller with Thompson Sampling for adaptive action selection (TSMPC), serving as a predictive-control baseline with analytic constraints and no learned uncertainty modeling.
@@ 4.2 Results
Fig. 3: Daily TIR by Patient Using TSODE
Figs. 2 and 3 illustrate the closed-loop performance of the proposed TSODE controller. Fig. 2 compares a representative 24-hour simulation against a conventional PID controller for the same virtual adult. While PID shows larger glucose excursions and periods of suboptimal regulation, TSODE maintains levels tightly within the target range of 70–180 mg/dL, achieving full time-in-range with no hypoglycemia. Fig. 3 extends this comparison across three adult subjects, where daily TIR values consistently cluster between 85% and 100%, reflecting stable glycemic control and strong generalization despite variability in metabolism and insulin sensitivity. Together, these results demonstrate that TSODE achieves safe, adaptive regulation superior to conventional controls.
Table 1 summarizes per-patient results and aggregate performance over 14 evaluation days. Across three representative adults and a six-subject cohort mean (adult#001, #002, #005, #008–#010), TSODE consistently achieves the highest TIR while limiting time below 70 mg/dL to under 10%, substantially outperforming PID and TSMPC, which incur up to 26% hypoglycemia time. These results demonstrate that TSODE maintains strong glycemic control and safety across both individual cases and heterogeneous patient dynamics. By combining meal-aware adaptation with NeuralODE forecasting, TSODE anticipates downward trends and caps unsafe doses, enhancing both performance and safety in autonomous insulin delivery.
Model Transferability and Generalization. To evaluate cross-patient generalization, the TSODE controller was trained on adults#001 and #002 and tested on an unseen subject (adult#005). As summarized in Table 1, the baseline achieved only 21.2% TIR, while the PID improved TIR to 75.1% but incurred 24.8% time below 70 mg/dL. In contrast, TSODE maintained a high TIR of 88.9% with only 3.8% time << 70 mg/dL, achieving comparable control with substantially reduced hypoglycemia risk. These results highlight the controller’s ability to generalize across unseen patients while preserving safety through NeuralODE-based probabilistic supervision.
@@ 5 Conclusion
We presented TSODE, a safety-aware controller that integrates Thompson Sampling reinforcement learning with NeuralODE-based probabilistic forecasting for automated insulin delivery. TSODE adaptively regulates glucose while certifying each insulin action through conformal safety checks and guardrail constraints. In the FDA-approved UVa/Padova simulator, TSODE achieved a mean time-in-range of 83.75% with only 8.83% time below 70 mg/dL, representing a 6–8 pp improvement in TIR and a 60% reduction in hypoglycemia compared to PID and TSMPC baselines. These results demonstrate that probabilistic safety filtering substantially enhances both stability and safety of closed-loop glucose control.